OpenStudy (anonymous):

Let f(x)=9-x^2, g(x)=3-x. Find (f-g)(x)

jimthompson5910 (jim_thompson5910):

(f-g)(x) = f(x) - g(x) ... by definition (f-g)(x) = (9-x^2) - (3-x)... plug in the given functions (f-g)(x) = 9-x^2 - 3 + x ... distribute I'll let you simplify